### v1

# VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning

Abstract
Vision-language navigation in unknown environments is crucial for mobile robots. In scenarios such as household assistance and rescue, mobile robots need to understand a human command, such as “find a person wearing black”. We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots. Unlike prior methods that rely on a single image-level feature similarity to guide a robot, we introduce the heuristic-vision-language (HVL) spatial reasoning for goal point selection. It combines pixel-wise vision-language features and heuristic exploration to enable efficient navigation to human-instructed instances in various environments robustly. We deploy VL-Nav on a four-wheel mobile robot and conduct comprehensive navigation tasks in various environments of different scales and semantic complexities, indoors and outdoors. Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation. Experimental results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and heuristic cues for mobile robots. (a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway. Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the nearby frontier point as the goal and successfully locating the missing person. The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points. (b) Apartment: The robot is tasked with “Go to the tall white trash bin.” It detected two different-sized white trash bins in bottom camera observation. However, it assigns a higher confidence score (0.98) to the taller bin than the shorter one (0.48). These pixel-wise VL features are incorporated into the spatial distribution to select the correct goal point, guiding the robot toward the taller bin.
I Introduction
Efficient navigation following human instructions in unseen environments is critical for autonomous robots, with various applications spanning from home assistance [28, 31, 32] to planetary exploration [25, 35, 27]. In such scenarios, robots are tasked with exploring and identifying targets instructed by humans, a challenge we define as vision-language navigation (VLN). Suppose a man wearing gray is reported missing and a command “find a man in gray” is given, the robot should not engage in random exploration. Instead, it should prioritize directions with visual features showing a stronger correlation to the language cues, such as detecting “gray cloth” within its field of view. Real-world VLN requires a navigation system that can: (1) interpret pixel-wise vision-language features, (2) adapt and perform robustly across various environments, and (3) operate in real-time on low-power platforms.
However, there are still no VLN systems that can fully address the three capabilities. Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [12]. Classical approaches [9, 10, 29, 33], while efficient, struggle to integrate vision-language features. End-to-end learning methods [17, 24, 7, 26, 22] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios. Modular learning approaches [3, 21, 13, 5, 4], on the other hand, demonstrate robust real-world performance [12]. However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning. The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches. For instance, Vision-Language Frontier Maps (VLFM) [36] utilized VLMs to extract language-grounded features directly from RGB images. This enables the creation of a semantic map, guiding exploration based on human semantic knowledge. However, their reliance on computationally intensive models limits their deployability on low-power platforms. Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained vision-language cues.
To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer. VL-Nav employs heuristic vision-language (HVL) spatial reasoning on both frontier-based and instance-based target points. VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection. It confines the search to a manageable field of view, thereby reducing computational overhead. Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate. To select the most informative goal point, we introduce HVL spatial reasoning. This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model. Each target point is then assigned a vision-language semantic score based on this distribution. Subsequently, a curiosity-driven heuristic weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into HVL scores. The target point with the highest HVL score is ultimately selected as the goal point. This HVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments. By combining pixel-wise vision-language features with heuristics via the novel HVL spatial reasoning, VL-Nav surpasses all the baselines, achieving human-like intelligent navigation while remaining computationally feasible for real-world deployment.
Our main contributions are as follows:
  • We propose VL-Nav, an efficient VLN system optimized for low-power robots, achieving robust real-time performance at 30 Hz with an onboard computer.
  • We introduce the heuristic-vision-language (HVL) spatial reasoning combining pixel-wise vision-language features and heuristic exploration for efficient and robust VLN..
  • We conduct comprehensive evaluations across four real-world settings, showing that VL-Nav outperforms prior methods by 44.15% in diverse environments.
II Related Works
Robot navigation methods can be broadly categorized into classical, end-to-end learning, and modular learning methods.
Classical Approaches:
Classical methods for navigation, including Simultaneous Localization and Mapping (SLAM)-based techniques, have been extensively studied over the last three decades [9, 10, 29, 33]. These methods typically build geometric maps of an environment using depth sensors or monocular RGB cameras [30, 19, 8, 23], while simultaneously localizing the robot relative to the growing map. Exploration is often guided by heuristic strategies like frontier-based methods [33, 11], while analytical planners are employed for obstacle avoidance and path planning. Although effective for traditional navigation tasks, these classical methods lack the ability to integrate vision-language features, making them insufficient for human-collaboration tasks like vision-language navigation.
End-to-End Learning:
End-to-end learning methods directly map sensory inputs to navigation actions using deep neural networks, typically trained with imitation learning (IL) or reinforcement learning (RL) losses [17, 24, 7, 26, 22]. Despite their ability to learn semantic priors for goal-directed exploration, end-to-end approaches suffer from significant limitations. These include high computational cost, reliance on large-scale training data, and a lack of interpretability. Furthermore, these methods often struggle to generalize beyond simulation due to domain gaps, as evidenced by significant performance drops when transitioning to real world [12].
Figure 2: An overview of the VL-Nav pipeline. VL-Nav processes inputs including prompts, RGB images, odometry poses, and LiDAR scans. The Vision-Language (VL) module conducts open-vocabulary pixel-wise detection to identify areas and objects related to the prompt, generating instance-based target points. Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map. Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool. VL-Nav employs heuristic-vision-language (HVL) spatial reasoning to select the most effective goal point from this pool for path planning.
Modular Learning:
Modular learning approaches seek to combine the advantages of classical and end-to-end methods by replacing specific components of the classical pipeline with learned modules. This modularity allows for the integration of semantic priors while maintaining system interpretability and efficiency [16, 18, 2, 1]. For example, semantic maps generated from RGB-D data have been used to guide navigation toward specific objects [3, 20, 21, 13]. Modular methods have shown promise in addressing Sim-to-Real transfer challenges by abstracting raw sensor data into higher-level representations, mitigating the impact of domain gaps [18, 1, 2]. However, existing modular approaches often rely heavily on real-world training data and struggle to utilize vision-language features.
Foundation Models for Navigation:
Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches. The Vision-Language Frontier Maps (VLFM) method [36] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration. VLFM achieves SOTA performance in the simulated environment, addressing the limitations of prior frontier-based and goal-oriented modular methods [33, 11, 3, 15]. However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN. Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for most real-world applications.
Strengths of VL-Nav:
VL-Nav significantly advances vision-language navigation (VLN) by addressing the limitations of classical, end-to-end, and modular learning methods. It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms. The integration of instance-based target points with heuristic vision-language (HVL) spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection. Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz. This makes VL-Nav not only practical for real-world applications but also superior in navigating complex environments efficiently.
III Methodology
Figure 3: A brief illustration of VL Scoring. The pixel-wise open-vocabulary detection results are transferred into the spatial distibution via the Gaussian mixture model regularized by the FOV weighting (the gray arrows in the figure). Then the frontier-based and the instance-based target points will be assigned with VL score based on the distribution Equation (3).
This section presents our VL-Nav pipeline (Fig. 2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and Heuristic-Vision-Language (HVL) spatial reasoning to select and publish a navigation goal. This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
III-A Rolling Occupancy Map: Occupancy Update
We represent the environment as a 2D occupancy grid $\mathcal{G}$, where each cell may be free $(0)$, unknown $(-1)$, or occupied $(100)$. Upon receiving new sensor data (a merged obstacle-and-terrain cloud $\mathcal{P}$), we update $\mathcal{G}$ using the following steps:
  1. 1.
    Expand if needed. If any newly observed points lie outside the current map boundaries, expand the grid in place to preserve historical data.
  2. 2.
    Clear stale obstacles in forward FOV. For each cell marked occupied within the robot’s forward field of view (FOV), check whether it still corresponds to a point in $\mathcal{P}$. Any “stale” obstacle cells are re-labeled as free.
  3. 3.
    Inflate new obstacles. For each point in $\mathcal{P}$ within the sensor range $R$, mark the corresponding cell as occupied, along with a local neighborhood determined by an inflation radius $r_{\mathrm{inflation}}$.
  4. 4.
    Raycast unknown cells to free. Perform line-of-sight raycasting from the robot’s pose $(x_{r},y_{r},\theta_{r})$ across its angular range $\theta\in[\theta_{r}-\tfrac{\mathrm{hfov}}{2},\;\theta_{r}+\tfrac{\mathrm{hfov}}{2}]$. Any unknown cells $(-1)$ intersected by a ray are marked free $(0)$.
Algorithm 1 summarizes these main steps, returning the updated grid $\mathcal{G}$. This occupancy update leverages consistent line-of-sight logic and inflation to support safe navigation.
Algorithm 1 Rolling Occupancy Map Update (Key Steps)
0:  Grid $\mathcal{G}$, merged point cloud $\mathcal{P}$, robot pose $(x_{r},y_{r},\theta_{r})$, inflation radius $r_{\mathrm{inflation}}$, half-FOV $\mathrm{hfov}/2$, sensor range $R$.
1:  $\textsc{ExpandIfNeeded}(\mathcal{G},\mathcal{P})$
[…]
5:  return  Updated grid $\mathcal{G}$
III-B Frontier-based Target Points
A cell $(m_{x},m_{y})$ is a frontier if it is free, $\mathcal{G}(m_{x},m_{y})=0$, and at least one neighbor is unknown, $\mathcal{G}(m_{x}+\Delta x,m_{y}+\Delta y)=-1$ for some $(\Delta x,\Delta y)$. After every grid update, only cells in the forward wedge are tested:
$\begin{split}\Bigl{|}\mathrm{Angle}\bigl{(}(m_{x},m_{y}),(x_{r},y_{r})\bigr{)}-\theta_{r}\Bigr{|}\;&\leq\;\frac{\mathrm{hfov}}{2},\\ \text{and}\quad\|(m_{x},m_{y})-(x_{r},y_{r})\|\;&\leq\;R.\end{split}$
A breadth-first search (BFS) clusters these frontier cells. Each cluster is represented by either a single centroid (for small clusters) or multiple sampled points (for large clusters) as the frontier-based target points.
III-C Instance-Based Target Points (IBTP)
A vision-language detector periodically reports candidate instance centers in the form $(q_{x},\;q_{y},\;\text{confidence})$, where $q_{x},q_{y}$ denote the estimated global coordinates for a potential target instance, and confidence quantifies how likely this detection is to match the desired instance. If confidence is larger than the detection threshold $\text{confidence is larger than the detection threshold }\tau_{\mathrm{det}}$, the candidate is retained; otherwise, it is discarded as too uncertain. Retained points will be downsampled via a voxel-grid filter if many candidates lie in close proximity.
This instance-based approach mimics human behavior when searching for an object: upon glimpsing something that might match the target, one naturally moves closer to confirm. VL-Nav likewise does not ignore intermediate detections. Instead, any candidate above a confidence threshold is treated as a valid goal candidate, allowing the robot to approach and verify whether it truly is an instance of interest. If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues, producing an accurate and robust strategy for zero-shot navigation toward a target instance.
III-D Heuristic-Vision-Language (HVL) Spatial Reasoning
Once frontier centroids and instance-based targets are gathered, the system computes an HVL score for each candidate goal $\mathbf{g}$. Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance. The angular offset is
$\Delta\theta\;=\;\mathrm{Angle}\bigl{(}\mathbf{g},(x_{r},y_{r})\bigr{)}\;-\;\theta_{r}.$
VL Score:
As shown in Fig. 3, a vision-language (VL) score $S_{\mathrm{VL}}(\mathbf{g})$ translates pixel-wise vision-language features into a Gaussian-mixture distribution over the robot’s horizontal field of view (FOV). Suppose the open-vocabulary detection model identifies $K$ likely directions or detection regions, each parameterized by $\bigl{(}\mu_{k},\sigma_{k},\alpha_{k}\bigr{)}$. Here, $\mu_{k}$ indicates the mean offset angle within the FOV, $\sigma_{k}$ encodes the detection’s angular uncertainty (in our implementation, a fixed constant of $0.1$ for all detections), and $\alpha_{k}$ is a confidence-based weight that captures how important that region is. By accumulating each region’s contribution in a mixture model, $S_{\mathrm{VL}}(\mathbf{g})$ naturally biases the robot toward goals in the direction of high-confidence detections that match the language prompt.
From a human perspective, this process resembles observing different areas that related to the target instance with varying levels of certainty. For instance, if someone is looking for a “red chair” in a large room, they might spot multiple red silhouettes at the right side of their vision, plus a clearer view of something red near the center. Although peripheral glimpses (“out of the corner of the eye”) often carry lower confidence, they are not ignored. Instead, we naturally merge these partial observations into a mental sense of “the chair is probably in one of these directions.” The system follows a similar approach, assigning each detection region a Gaussian shape around $\mu_{k}$, scaled by $\sigma_{k}$ and weighted by $\alpha_{k}$. Inspired by VLFM [36], we then multiply by $C(\Delta\theta)$, a “view confidence” term that downweights detections with large angular offsets from the central field of view. The VL score is computed as:
$S_{\mathrm{VL}}(\mathbf{g})\;=\;\sum_{k=1}^{K}\alpha_{k}\;\exp\!\Bigl{(}-\tfrac{1}{2}\Bigl{(}\tfrac{\Delta\theta-\mu_{k}}{\sigma_{k}}\Bigr{)}^{2}\Bigr{)}\;\cdot\;C\bigl{(}\Delta\theta\bigr{)},$
$C(\Delta\theta)=\cos^{2}\!\biggl{(}\frac{\Delta\theta}{(\theta_{\mathrm{fov}}/2)}\cdot\frac{\pi}{2}\biggr{)}.$
Heuristic Cues:
We add two heuristic terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
(1) Distance Weighting. We define
$S_{\mathrm{dist}}(\mathbf{g})\;=\;\frac{1}{1+d(\mathbf{x}_{r},\mathbf{g})},$
so that nearer goals receive slightly higher scores. This factor is especially important on a real robot, as shorter travel distances can significantly reduce energy consumption while preventing needless wandering. Although the distance term alone does not guarantee an optimal path, it helps prioritize goals that can be reached sooner.
(2) Unknown-Area Weighting. We further encourage curiosity-driven exploration by measuring how many unknown cells lie around $\mathbf{g}$. First, we define the ratio of unknown as:
[…]
where a local BFS radiates outward from $\mathbf{g}$, counting how many cells are unknown versus how many are reachable (i.e., not blocked by obstacles). Larger $\mathrm{ratio}(\mathbf{g})$ implies that moving to $\mathbf{g}$ may reveal significant unknown space and get more information. To translate this raw ratio into a normalized weighting score, we apply an exponential mapping:
[…]
where $k$ is an adjustable parameter that controls how rapidly the score increases from 0 toward 1.
Combined HVL Score:
The final scores of frontier-based goals combine these three components:
$S_{\mathrm{HVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights. Note that the VL score is assigned when a point is initially detected, whereas the heuristic terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated. For instance-based goals, these heuristic cues are omitted because their primary purpose is verification rather than exploration.
III-E Select & Publish Goal Point
After evaluating HVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate. Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
Algorithm 2 Select and Publish Goal Point
0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $HVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
1:  $\textit{bestScore}\leftarrow-\infty$
[…]
3:  for all instances $\mathbf{o}_{i}$ do
4:     if $\|\mathbf{o}_{i}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
[…]
6:        if $s>\textit{bestScore}$ then
7:           $\textit{bestScore}\leftarrow s$
8:           $\textit{bestGoal}\leftarrow\mathbf{o}_{i}$
9:        end if
10:     end if
11:  end for
[…]
13:     for all frontiers $\mathbf{f}_{j}$ do
14:        if $\|\mathbf{f}_{j}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
[…]
16:           if $s>\textit{bestScore}$ then
17:              $\textit{bestScore}\leftarrow s$
18:              $\textit{bestGoal}\leftarrow\mathbf{f}_{j}$
19:           end if
20:        end if
21:     end for
22:  end if
[…]
24:     Publish bestGoal to /selected_goal_point
25:  end if
26:  return
If no candidate exceeds the distance threshold $\delta_{\mathrm{reached}}$, the robot withholds any new goal. Once a goal is published, a global planner updates obstacle polygons and replans a safe route. A local planner refines each global waypoint into short-horizon commands for real-time obstacle avoidance, supporting incremental exploration in unknown environments.
III-F Path Planning with FAR Planner
Once an HVL goal is selected, we hand off point-goal path planning to the FAR Planner [34], which represents obstacles polygonally and updates a visibility graph in real time. This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods. A local planner refines FAR Planner’s waypoints into short-horizon velocity commands, ensuring swift reactions to new obstacles and seamless integration with VL-Nav’s zero-shot vision-language objectives.
IV Experiments
Figure 4: Four different real-world experiment environments.
(a) Office
(b) Apartment
(c) Outdoor
(d) Hallway
Figure 5: Top-down view of the trajectories comparison on the value maps with the detection results across the four different environments.
TABLE I: Vision-language navigation performance in 4 unseen environments (SR and SPL).
Method SR (%) SPL
Hallway Office Apartment Outdoor Hallway Office Apartment Outdoor
Frontier Exploration 40.0 41.7 55.6 33.3 0.239 0.317 0.363 0.189
VLFM [36] 53.3 75.0 66.7 44.4 0.366 0.556 0.412 0.308
VL-Nav w/o IBTP 66.7 83.3 70.2 55.6 0.593 0.738 0.615 0.573
VL-Nav w/o Heuristics 73.3 86.3 66.7 55.6 0.612 0.743 0.631 0.498
VL-Nav 86.7 91.7 88.9 77.8 0.672 0.812 0.733 0.637
IV-A Experimental Setting
We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [36], (3) VLNav without instance-based target points, (4) VLNav without heuristic terms, and (5) the full VLNav configuration. Because the original VLFM relies on BLIP-2 [14], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM. Each method is tested under the same conditions to ensure a fair comparison of performance.
Environments:
We consider four distinct environments (shown in Fig. 4), each with a specific combination of semantic complexity (High, Medium, or Low) and size (Big, Mid, or Small). Concretely, we use a Hallway (Medium & Big), an Office (High & Mid), an Outdoor area (Low & Big), and an Apartment (High & Small). In each environment, we evaluate five methods using three language prompts, yielding a diverse range of spatial layouts and semantic challenges. This setup provides a rigorous assessment of each method’s adaptability.
Language-Described Instance:
We define nine distinct, uncommon human-described instances to serve as target objects or persons during navigation. Examples include phrases such as “tall white trash bin,” “there seems to be a man in white,” “find a man in gray,” “there seems to be a black chair,” “tall white board,” and “there seems to be a fold chair.” The variety in these descriptions ensures that the robot must rely on vision-language understanding to accurately locate these targets.
Robots and Sensor Setup: All experiments are conducted using a four-wheel Rover equipped with a Livox Mid-360 LiDAR. The LiDAR is tilted by approximately 23 degrees to the front to achieve a ±$\pm 30$ degrees vertical FOV coverage closely aligned with the forward camera’s view. An Intel RealSense D455 RGB-D camera, tilted upward by 7 degrees to detect taller objects, provides visual observation, though its depth data are not used for positioning or mapping. LiDAR measurements are a primary source of mapping and localization due to their higher accuracy. The whole VL-Nav system runs on an NVIDIA Jetson Orin NX on-board computer.
IV-B Main Results
We validate the proposed VL-Nav system in real-robot experiments across four distinct environments (Hallway, Office, Apartment, and Outdoor), each featuring different semantic levels and sizes. Building on the motivation articulated in Section I, we focus on evaluating VL-Nav’s ability to (1) interpret fine-grained vision-language features and conduct robust VLN, (2) explore efficiently in unfamiliar spaces across various environments, and (3) run in real-time on resource-constrained platforms. Fig. 5 presents a top-down comparison of trajectories and detection results on the value map.
Figure 6: Plots of performance in different environments sizes and semantic comlexities.
Overall Performance:
As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments. In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating HVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
Effect of Instance-Based Target Points (IBTP):
We note a marked improvement when enabling IBTP: the variant without IBTP lags behind, particularly in complex domains like the Apartment and Office. As discussed in Section III, IBTP allows VL-Nav to pursue and verify tentative detections with confidence above a threshold, mirroring human search behavior. This pragmatic mechanism prevents ignoring possible matches to the target description and reduces overall travel distance to confirm or discard candidate objects.
Heuristics Contributions:
The Heuristic Score is also significant to VL-Nav’s performance. It merges two key components:
  • Distance Weighting: Preventing easily select very far way goals to reduce travel time and energy consumption which is extremely important for the efficiency (metrics SPL) in the large-size environments.
  • Unknown-Area Heuristic: Rewards navigation toward regions that yield more information.
Our ablations reveal that removing the distance-scoring element (VL-Nav w/o heuristics) degrades both SR and SPL, particularly in the more cluttered environments. Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of HVL addresses a complementary aspect of semantic navigation.
Comparison to VLFM:
Although the VLFM approach [36] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and HVL-based spatial reasoning. Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the heuristic cues via the HVL spatial reasoning. These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our HVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
Summary of Findings:
In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments. By fusing frontier-based target points detection, instance-based target points, and the HVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency. The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
V Discussion
V-A Sizes of Environments vs. SR and SPL
Our experiments grouped environments by size into three categories: Small (Apartment), Mid (Office), and Big (averaging Hallway and Outdoor). As shown in Table I and the line plots in Fig. 6, classical Frontier Exploration suffers a marked performance decline in larger spaces, frequently exceeding the path length limit. Its success rate drops from about 55.6% in the Small environment to only 36.7% in the Big, and its SPL falls from 0.564 to 0.214. This illustrates how naive 360° frontier-based approaches grow increasingly inefficient in larger, more open environments.
By contrast, VL-Nav remains robust across all sizes. Although there is a slight decrease when moving from Mid (91.7% SR, 0.852 SPL) to Big (82.3% SR, 0.677 SPL), its overall performance remains substantially higher than the baselines. A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o Heuristic), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
These findings underscore two key factors. First, HVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig. 1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments. Second, instance-based target points (IBTP) enable more targeted navigation, sustaining high success rates even as map size grows.
V-B Semantic Complexities of Environments vs. SR and SPL
We also vary the environments by semantic complexity, classifying them as Low (Outdoor), Medium (Hallway), and High (averaging Office and Apartment). An interesting trend emerges: all methods show better performance in more semantically rich environments. The likely explanation is that structured indoor spaces (e.g., with walls or furniture) provide stronger cues for the detection and segmentation models used by all five methods, leading to more accurate navigation.
These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context. Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and heuristic exploration. The fusion of the HVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
V-C Dynamic Environment Handling
Handling dynamic environments is crucial for robust navigation in real-world scenarios, where changes such as moving obstacles and varying terrain can impact robot performance. VL-Nav incorporates several strategies to dynamically adapt to environmental changes, ensuring reliability and efficiency.
To handle the dynamic environment, we maintain a dynamic occupancy grid map that integrates dynamic terrain analysis and path planning within the FAR planner [34]. This allows the system to continuously update its understanding of the environment and adjust navigation strategies accordingly. The dynamic grid map ensures that real-time changes, such as shifting obstacles and terrain conditions, are accurately reflected, providing a more robust foundation for dynamic vision-language navigation in the social environment.
V-D Real-Robot Efficiency Optimization
Although VL-Nav is designed as a general navigation framework, we introduce several optimizations to ensure real-world feasibility and high performance on mobile robots.
Visual-Language Module Variants:
The first optimization involves selecting an efficient model variant for the visual-language (or object-detection) pipeline. In our tests, we ran multiple versions of YOLO-World [6] detectors with different input dimensions and TensorRT precision modes. For instance, one configuration used a $256\times 320$ input size in a standard GPU runtime, another used TensorRT (FP32) at the same resolution, and a third tested TensorRT (FP16) at $480\times 640$. Empirically, the original $256\times 320$ configuration (run without TensorRT) achieved about $23\%$ higher success rates than higher-resolution FP16 modes. This suggests that moderate input resolutions can strike a more favorable balance between detection accuracy and real-time performance, which is crucial for robust and efficient continuous on-robot inference.
Rolling Occupancy Grid vs. Fixed Large Grid:
A second optimization lies in how the occupancy grid is maintained. Rather than keeping a large, fixed-size global grid (common in many simulation environments), we implement a rolling grid on the real robot, dynamically shifting or expanding it only as new sensor data demand. This rolling approach keeps memory usage proportional to the active exploration area, lowering both storage overhead and BFS/cluster computations for partial frontier detection. Moreover, since real-world robots typically explore incrementally, discarding far-off map regions has minimal impact on decision-making.
Additional System-Level Improvements:
Finally, we incorporate various practical measures to ensure robust navigation under real-world constraints. For instance, the robot’s local planner and velocity commands are monitored to detect “stuck” situations; if the robot remains motionless beyond a time threshold, the system discards its current goal and reselects a new one. Similarly, sensor frequency and map update rates are tuned so that occupancy-grid expansions do not saturate available CPU/GPU resources. These measures, alongside the rolling-grid and efficient detection models, help keep VL-Nav lightweight and reactive.
VI Limitation and Conclusion
VI-A Limitations
Despite the promising results demonstrated by VL-Nav, several limitations remain. One major challenge is the system’s ability to handle complex language descriptions. It struggles with spatial descriptions that contain hidden object references and objects described with specific textual annotations, which can affect navigation accuracy. Another limitation lies in the reliance on manually defined thresholds for various navigation conditions like the lightning condition. These thresholds may not generalize well across different environments and scenarios, requiring further investigation into adaptive or learning-based approaches to threshold tuning.
VI-B Conclusion
In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms. By combining pixel-wise vision-language features with a heuristic exploration strategy, our HVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings. Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%. The key to these gains lies in effectively leveraging semantic cues from vision-language embeddings to prioritize frontiers and potential object instances, thereby aligning navigation decisions more closely with human-like reasoning.
Future work will explore extending VL-Nav to handle more complex instructions that involve multi-step tasks, temporal reasoning (e.g., tracking a moving target), and dynamic environments. Additionally, further integration with large language models could enable more nuanced command parsing and open-vocabulary object detection for even broader applicability. We believe these directions will help advance the capability of VLN systems and bring us closer to deploying versatile, robust robotic assistants in real-world settings.
Acknowledgments
This work was partially supported by the Sony Faculty Innovation Award and the DARPA grant DARPA-PS-23-13. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DARPA.

### v2

@@
 We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots.
-Unlike prior methods that rely on a single image-level feature similarity to guide a robot, we introduce the heuristic-vision-language (HVL) spatial reasoning for goal point selection.
-It combines pixel-wise vision-language features and heuristic exploration to enable efficient navigation to human-instructed instances in various environments robustly.
-We deploy VL-Nav on a four-wheel mobile robot and conduct comprehensive navigation tasks in various environments of different scales and semantic complexities, indoors and outdoors.
+Unlike prior methods that rely on a single image-level feature similarity to guide a robot, our method integrates pixel-wise vision-language features with curiosity-driven exploration.
+This approach enables robust navigation to human-instructed instances across diverse environments.
+We deploy VL-Nav on a four-wheel mobile robot and evaluate its performance through comprehensive navigation tasks in both indoor and outdoor environments, spanning different scales and semantic complexities.
 Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation.
-Experimental results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
-Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and heuristic cues for mobile robots.
+Results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
+Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and curiosity-based exploration for mobile robots.
 (a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway.
-Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the nearby frontier point as the goal and successfully locating the missing person.
+Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the most VL-correlated goal point and successfully locating the missing person.
 The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points.
@@
 To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer.
-VL-Nav employs heuristic vision-language (HVL) spatial reasoning on both frontier-based and instance-based target points.
+VL-Nav employs the spatial reasoning on both frontier-based and instance-based target points.
 VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection.
@@
 Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate.
-To select the most informative goal point, we introduce HVL spatial reasoning.
+To select the most informative goal point, we introduce CVL spatial reasoning.
 This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model.
 Each target point is then assigned a vision-language semantic score based on this distribution.
-Subsequently, a curiosity-driven heuristic weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into HVL scores.
-The target point with the highest HVL score is ultimately selected as the goal point.
-This HVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
+Subsequently, a curiosity-driven weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into CVL scores.
+The target point with the highest CVL score is ultimately selected as the goal point.
+This CVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
 Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments.
-By combining pixel-wise vision-language features with heuristics via the novel HVL spatial reasoning, VL-Nav surpasses all the baselines, achieving human-like intelligent navigation while remaining computationally feasible for real-world deployment.
+By combining pixel-wise vision-language features with curiosity-based exploration via the novel CVL spatial reasoning, VL-Nav surpasses all the baselines, achieving intelligent navigation while remaining computationally feasible for real-world deployment.
 Our main contributions are as follows:
   • We propose VL-Nav, an efficient VLN system optimized for low-power robots, achieving robust real-time performance at 30 Hz with an onboard computer.
-  • We introduce the heuristic-vision-language (HVL) spatial reasoning combining pixel-wise vision-language features and heuristic exploration for efficient and robust VLN..
+  • We empower VL-Nav with spatial reasoning by integrating pixel-wise vision-language features and curiosity-driven exploration for more efficient VLN..
   • We conduct comprehensive evaluations across four real-world settings, showing that VL-Nav outperforms prior methods by 44.15% in diverse environments.
@@
 Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool.
-VL-Nav employs heuristic-vision-language (HVL) spatial reasoning to select the most effective goal point from this pool for path planning.
+VL-Nav employs spatial reasoning to select the most effective goal point from this pool for path planning.
 Modular Learning:
@@
 It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms.
-The integration of instance-based target points with heuristic vision-language (HVL) spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection.
+The integration of instance-based target points with spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection.
 Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz.
-This makes VL-Nav not only practical for real-world applications but also superior in navigating complex environments efficiently.
+This makes VL-Nav not only practical for real-world applications but also superior in navigating complex and various environments efficiently.
 III Methodology
@@
 This section presents our VL-Nav pipeline (Fig.
-2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and Heuristic-Vision-Language (HVL) spatial reasoning to select and publish a navigation goal.
+2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and curiosity-vision-language (CVL) Scoring Policy to select and publish a navigation goal.
 This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
@@
 If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues, producing an accurate and robust strategy for zero-shot navigation toward a target instance.
-III-D Heuristic-Vision-Language (HVL) Spatial Reasoning
-Once frontier centroids and instance-based targets are gathered, the system computes an HVL score for each candidate goal $\mathbf{g}$.
+III-D CVL Scoring Policy
+Once frontier centroids and instance-based targets are gathered, the system computes an CVL score for each candidate goal $\mathbf{g}$.
 Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance.
@@
 $C(\Delta\theta)=\cos^{2}\!\biggl{(}\frac{\Delta\theta}{(\theta_{\mathrm{fov}}/2)}\cdot\frac{\pi}{2}\biggr{)}.$
-Heuristic Cues:
-We add two heuristic terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
+Curiosity Cues:
+We add two curiosity terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
 (1) Distance Weighting. We define
@@
 where $k$ is an adjustable parameter that controls how rapidly the score increases from 0 toward 1.
-Combined HVL Score:
+Combined CVL Score:
 The final scores of frontier-based goals combine these three components:
-$S_{\mathrm{HVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
+$S_{\mathrm{CVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
 where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights.
-Note that the VL score is assigned when a point is initially detected, whereas the heuristic terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated.
-For instance-based goals, these heuristic cues are omitted because their primary purpose is verification rather than exploration.
+Note that the VL score is assigned when a point is initially detected, whereas the curiosity terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated.
+For instance-based goals, these curiosity cues are omitted because their primary purpose is verification rather than exploration.
 III-E Select & Publish Goal Point
-After evaluating HVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate.
+After evaluating CVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate.
 Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
 Algorithm 2 Select and Publish Goal Point
-0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $HVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
+0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $CVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
 1:  $\textit{bestScore}\leftarrow-\infty$
@@
 III-F Path Planning with FAR Planner
-Once an HVL goal is selected, we hand off point-goal path planning to the FAR Planner [34], which represents obstacles polygonally and updates a visibility graph in real time.
+Once an CVL goal is selected, we hand off point-goal path planning to the FAR Planner [34], which represents obstacles polygonally and updates a visibility graph in real time.
 This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods.
@@
 VL-Nav w/o IBTP 66.7 83.3 70.2 55.6 0.593 0.738 0.615 0.573
-VL-Nav w/o Heuristics 73.3 86.3 66.7 55.6 0.612 0.743 0.631 0.498
+VL-Nav w/o curiosity 73.3 86.3 66.7 55.6 0.612 0.743 0.631 0.498
 VL-Nav 86.7 91.7 88.9 77.8 0.672 0.812 0.733 0.637
 IV-A Experimental Setting
-We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [36], (3) VLNav without instance-based target points, (4) VLNav without heuristic terms, and (5) the full VLNav configuration.
+We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [36], (3) VLNav without instance-based target points, (4) VLNav without curiosity terms, and (5) the full VLNav configuration.
 Because the original VLFM relies on BLIP-2 [14], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM.
@@
 As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments.
-In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating HVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
+In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating CVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
 Effect of Instance-Based Target Points (IBTP):
@@
 This pragmatic mechanism prevents ignoring possible matches to the target description and reduces overall travel distance to confirm or discard candidate objects.
-Heuristics Contributions:
-The Heuristic Score is also significant to VL-Nav’s performance. It merges two key components:
+Curiosity Contributions:
+The curiosity Score is also significant to VL-Nav’s performance. It merges two key components:
   • Distance Weighting: Preventing easily select very far way goals to reduce travel time and energy consumption which is extremely important for the efficiency (metrics SPL) in the large-size environments.
-  • Unknown-Area Heuristic: Rewards navigation toward regions that yield more information.
-Our ablations reveal that removing the distance-scoring element (VL-Nav w/o heuristics) degrades both SR and SPL, particularly in the more cluttered environments.
-Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of HVL addresses a complementary aspect of semantic navigation.
+  • Unknown-Area Weighting: Rewards navigation toward regions that yield more information.
+Our ablations reveal that removing the distance-scoring element (VL-Nav w/o curiosity) degrades both SR and SPL, particularly in the more cluttered environments.
+Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of CVL addresses a complementary aspect of semantic navigation.
 Comparison to VLFM:
-Although the VLFM approach [36] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and HVL-based spatial reasoning.
-Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the heuristic cues via the HVL spatial reasoning.
-These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our HVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
+Although the VLFM approach [36] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and CVL-based spatial reasoning.
+Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the curiosity cues via the CVL spatial reasoning.
+These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our CVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
 Summary of Findings:
 In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments.
-By fusing frontier-based target points detection, instance-based target points, and the HVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency.
+By fusing frontier-based target points detection, instance-based target points, and the CVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency.
 The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
@@
 By contrast, VL-Nav remains robust across all sizes.
-Although there is a slight decrease when moving from Mid (91.7% SR, 0.852 SPL) to Big (82.3% SR, 0.677 SPL), its overall performance remains substantially higher than the baselines.
-A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o Heuristic), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
+Although there is a slight decrease when moving from Mid (91.7% SR, 0.812 SPL) to Big (82.3% SR, 0.655 SPL), its overall performance remains substantially higher than the baselines.
+A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o curiosity), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
 These findings underscore two key factors.
-First, HVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig.
+First, CVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig.
 1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments.
@@
 These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context.
-Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and heuristic exploration.
-The fusion of the HVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
+Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and curiosity-based exploration .
+The fusion of the CVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
 V-C Dynamic Environment Handling
@@
 In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms.
-By combining pixel-wise vision-language features with a heuristic exploration strategy, our HVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings.
+By combining pixel-wise vision-language features with a curiosity-based exploration strategy, our CVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings.
 Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%.

### v3

@@
 Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [13].
-Classical approaches [9, 10, 29, 33], while efficient, struggle to integrate vision-language features.
+Classical approaches [9, 10, 31, 35], while efficient, struggle to integrate vision-language (VL) features.
 End-to-end learning methods [19, 26, 7, 28, 24] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios.
@@
 However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning.
-The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches.
+The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches [12, 38, 16, 39].
 For instance, Vision-Language Frontier Maps (VLFM) [39] utilized VLMs to extract language-grounded features directly from RGB images.
@@
 However, their reliance on computationally intensive models limits their deployability on low-power platforms.
-Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained vision-language cues.
+Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained VL cues.
 To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer.
@@
 Foundation Models for Navigation:
-Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches.
+Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches [12, 38, 16, 39].
 The Vision-Language Frontier Maps (VLFM) method [39] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration.
@@
 However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN.
-Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for most real-world applications.
+Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for mobile-robot applications.
 Strengths of VL-Nav:

### v4

@@
-VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning
+VL-Nav: A Neuro-Symbolic Approach for Reasoning-based Vision-Language Navigation
 Abstract
-Vision-language navigation in unknown environments is crucial for mobile robots.
-In scenarios such as household assistance and rescue, mobile robots need to understand a human command, such as “find a person wearing black”.
-We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots.
-Unlike prior methods that rely on a single image-level feature similarity to guide a robot, our method integrates pixel-wise vision-language features with curiosity-driven exploration.
-This approach enables robust navigation to human-instructed instances across diverse environments.
-We deploy VL-Nav on a four-wheel mobile robot and evaluate its performance through comprehensive navigation tasks in both indoor and outdoor environments, spanning different scales and semantic complexities.
-Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation.
-Results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
-Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and curiosity-based exploration for mobile robots.
-(a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway.
-Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the most VL-correlated goal point and successfully locating the missing person.
-The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points.
-(b) Apartment: The robot is tasked with “Go to the tall white trash bin.” It detected two different-sized white trash bins in bottom camera observation.
-However, it assigns a higher confidence score (0.98) to the taller bin than the shorter one (0.48).
-These pixel-wise VL features are incorporated into the spatial distribution to select the correct goal point, guiding the robot toward the taller bin.
+Navigating unseen, large-scale environments based on complex and abstract human instructions remains a formidable challenge for autonomous mobile robots.
+Addressing this requires robots to infer implicit semantics and efficiently explore large-scale task spaces.
+However, existing methods, ranging from end-to-end learning to foundation model-based modular architectures, often lack the capability to decompose complex tasks or employ efficient exploration strategies, leading to robot aimless wandering or target recognition failures.
+To address these limitations, we propose VL-Nav, a neuro-symbolic (NeSy) vision-language navigation system.
+The proposed system intertwines neural reasoning with symbolic guidance through two core components: (1) a NeSy task planner that leverages a symbolic 3D scene graph and image memory system to enhance the vision-language models’ (VLMs) neural reasoning capabilities for task decomposition and replanning; and (2) a NeSy exploration system that couples neural semantic cues with the symbolic heuristic function to efficiently gather the task-related information while minimizing unnecessary repeat travel during exploration.
+Validated on the DARPA TIAMAT Challenge navigation tasks, our system achieved an 83.4% success rate (SR) in indoor environments and 75% in outdoor scenarios.
+VL-Nav achieved an 86.3% SR in real-world experiments, including a challenging 483-meter run.
+Finally, we validate the system with complex instructions in a 3D multi-floor scenario.
+Project Website: https://sairlab.org/vlnav/
 I Introduction
-Efficient navigation following human instructions in unseen environments is critical for autonomous robots, with various applications spanning from home assistance [30, 33, 34] to planetary exploration [27, 37, 29].
-In such scenarios, robots are tasked with exploring and identifying targets instructed by humans, a challenge we define as vision-language navigation (VLN).
-Suppose a man wearing gray is reported missing and a command “find a man in gray” is given, the robot should not engage in random exploration.
-Instead, it should prioritize directions with visual features showing a stronger correlation to the language cues, such as detecting “gray cloth” within its field of view.
-Real-world VLN requires a navigation system that can: (1) interpret pixel-wise vision-language features, (2) adapt and perform robustly across various environments, and (3) operate in real-time on low-power platforms.
-However, there are still no VLN systems that can fully address the three capabilities.
-Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [13].
-Classical approaches [9, 10, 31, 35], while efficient, struggle to integrate vision-language (VL) features.
-End-to-end learning methods [19, 26, 7, 28, 24] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios.
-Modular learning approaches [3, 23, 14, 5, 4], on the other hand, demonstrate robust real-world performance [13].
-However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning.
-The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches [12, 38, 16, 39].
-For instance, Vision-Language Frontier Maps (VLFM) [39] utilized VLMs to extract language-grounded features directly from RGB images.
-This enables the creation of a semantic map, guiding exploration based on human semantic knowledge.
-However, their reliance on computationally intensive models limits their deployability on low-power platforms.
-Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained VL cues.
-To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer.
-VL-Nav employs the spatial reasoning on both frontier-based and instance-based target points.
-VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection.
-It confines the search to a manageable field of view, thereby reducing computational overhead.
-Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate.
-To select the most informative goal point, we introduce CVL spatial reasoning.
-This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model.
-Each target point is then assigned a vision-language semantic score based on this distribution.
-Subsequently, a curiosity-driven weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into CVL scores.
-The target point with the highest CVL score is ultimately selected as the goal point.
-This CVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
-Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments.
-By combining pixel-wise vision-language features with curiosity-based exploration via the novel CVL spatial reasoning, VL-Nav surpasses all the baselines, achieving intelligent navigation while remaining computationally feasible for real-world deployment.
+Efficient navigation following complex human instructions in completely unseen environments remains a formidable challenge for autonomous mobile robots.
+For example, consider a navigation task in the DARPA TIAMAT challenge [22]: “Today’s weather report indicates rain.
+Help Rob find an umbrella, an appropriate jacket, and shoes.” This marks a critical shift from naive command following vision-and-language navigaion [14] to reasoning-based vision-language navigation (VLN), where the robot must bridge massive logical gaps.
+The robot must first accurately understand the implicit semantics of the abstract instruction, reasoning out semantic inferences (e.g., “rain” implies waterproof gear) and performing object disambiguation to identify that the required targets are an umbrella, a rain jacket, and rain boots.
+This is essential to prevent the robot from looking for a normal jacket or sneakers as the target objects.
+Subsequently, it must efficiently explore the unknown environment to find all these required items within a limited task time.
+Figure 1: Given the complex instruction, VL-Nav autonomously navigates a university building.
+Top: The generated 3D global map and robot trajectory (green line).
+The projected value map (rainbow squares on the floor) visualizes the NeSy scoring policy, which prioritizes task-relevant areas to guide efficient exploration.
+Bottom: Key reasoning milestones where the system (1) identifies stairs for floor transition, (2) grounds the target object using spatial relation (on black box), and (3) verifies the recipient based on visual attributes.
+However, a significant gap remains in addressing such complex VLN.
+Classical semantic-free approaches [9, 10, 33] lack the linguistic reasoning necessary to comprehend abstract human instructions.
+Conversely, end-to-end learning methods, such as RL-based [19, 28, 7] and VLA-based approaches [38, 5, 37], are notoriously data-hungry and struggle with sim-to-real transfer.
+While recent foundation model-based modular architectures [36, 12, 35, 39] integrate advanced VLMs for zero-shot single-object navigation, their tight coupling of target verification with exploration constrains inferential capabilities, frequently resulting in the identification of incorrect items (e.g., random shoes instead of rain boots).
+Furthermore, most of these methods over-rely on neural semantic cues while neglecting the potential of geometric frontiers, leading to aimless wandering within explored regions.
+Consequently, existing approaches fundamentally lack two critical capabilities: (1) Accurate Task Reasoning to parse and decompose abstract multi-target instructions; and (2) Efficient Exploration Strategies to rapidly locate multiple targets in large-scale environments.
+To seamlessly bridge this gap, we propose a NeSy vision-language navigation system that fundamentally intertwines neural reasoning with symbolic guidance.
+Specifically, our VL-Nav architecture consists of two main modules: (1) NeSy Task Planner: This planner logically decomposes abstract, multi-target instructions into atomic subtasks, dynamically issuing either exploration subtasks to gather information or go to subtasks to navigate to and report target objects.
+Crucially, the planner grounds VLM’s neural reasoning in a unified symbolic memory comprising a 3D scene graph and object-centric image memory, anchoring VLM decisions in precise spatial awareness and historical context.
+(2) NeSy Exploration System: Serving as the exploration backbone, this module couples neural semantic cues from lightweight VLMs with a symbolic heuristic function.
+This integration rapidly guides the robot through unknown environments to discover targets while minimizing unnecessary travel.
+We comprehensively validate our proposed VL-Nav system in both simulations and extensive real-world robot environments.
+When evaluated in the navigation tasks in the DARPA TIAMAT challenge, our system demonstrated remarkable performance, achieving an 83.4% success rate (SR) in indoor settings and 75% in outdoor environments.
+Furthermore, real-world deployment across four diverse environments confirms the capability of generalizing from highly occluded, indoor domains to unstructured outdoor environments.
+Our real-world experiments achieved an 86.3% success rate, covering challenging scenarios such as a 483-meter long-range trajectory.
+Additionally, we showcased a complex multi-floor demonstration as shown in Fig.
+1.
 Our main contributions are as follows:
-  • We propose VL-Nav, an efficient VLN system optimized for low-power robots, achieving robust real-time performance at 30 Hz with an onboard computer.
-  • We empower VL-Nav with spatial reasoning by integrating pixel-wise vision-language features and curiosity-driven exploration for more efficient VLN..
-  • We conduct comprehensive evaluations across four real-world settings, showing that VL-Nav outperforms prior methods by 44.15% in diverse environments.
+  • We introduce VL-Nav, a NeSy system designed to resolve reasoning-based VLN tasks by intertwining neural semantic understanding with symbolic guidance.
+  • We introduce a robust NeSy Task Planner to guide VLM reasoning and replanning within a unified symbolic memory.
+When coupled with our NeSy Exploration System, this approach facilitates rapid multi-target discovery and minimizes unnecessary travel.
+  • VL-Nav achieves 83.4% and 75% SR in the indoor and outdoor settings on the navigation tasks of the DARPA TIAMAT Challenge, respectively.
+In real-world experiments, it attains an 86.3% SR, including long-range runs of up to 483 meters and demonstrations of complex tasks in multi-floor environments.
 II Related Works
@@
 End-to-End Learning:
-End-to-end learning methods directly map sensory inputs to navigation actions using deep neural networks, typically trained with imitation learning (IL) or reinforcement learning (RL) losses [19, 26, 7, 28, 24].
-Despite their ability to learn semantic priors for goal-directed exploration, end-to-end approaches suffer from significant limitations.
-These include high computational cost, reliance on large-scale training data, and a lack of interpretability.
-Furthermore, these methods often struggle to generalize beyond simulation due to domain gaps, as evidenced by significant performance drops when transitioning to real world [13].
-Figure 2: An overview of the VL-Nav pipeline.
-VL-Nav processes inputs including prompts, RGB images, odometry poses, and LiDAR scans.
-The Vision-Language (VL) module conducts open-vocabulary pixel-wise detection to identify areas and objects related to the prompt, generating instance-based target points.
-Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map.
-Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool.
-VL-Nav employs spatial reasoning to select the most effective goal point from this pool for path planning.
+End-to-end learning methods seek to directly map sensory inputs and instructions to navigation actions.
+Traditionally, these approaches relied on deep neural networks trained via imitation learning (IL) or reinforcement learning (RL) [19, 28, 7, 30, 26].
+More recently, Vision-Language-Action (VLA) models [38, 5, 37] have emerged, attempting to unify vision, language, and continuous control into single massive transformer architectures.
+Despite their ability to inherently learn semantic priors for goal-directed exploration, these end-to-end approaches suffer from significant limitations when applied to real-world, multi-target VLN.
+They are notoriously data-hungry, computationally expensive to run in real-time, and lack interpretability.
+Most critically, they frequently fail to generalize beyond their training distributions, leading to severe performance degradation during sim2real transfer [13].
 Modular Learning:
@@
 However, existing modular approaches often rely heavily on real-world training data and struggle to utilize vision-language features.
-Foundation Models for Navigation:
-Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches [12, 38, 16, 39].
-The Vision-Language Frontier Maps (VLFM) method [39] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration.
-VLFM achieves SOTA performance in the simulated environment, addressing the limitations of prior frontier-based and goal-oriented modular methods [35, 11, 3, 17].
-However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN.
-Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for mobile-robot applications.
-Strengths of VL-Nav:
-VL-Nav significantly advances vision-language navigation (VLN) by addressing the limitations of classical, end-to-end, and modular learning methods.
-It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms.
-The integration of instance-based target points with spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection.
-Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz.
-This makes VL-Nav not only practical for real-world applications but also superior in navigating complex and various environments efficiently.
+Foundation Model-Based Modular Learning Methods:
+Recent advances in foundation models like large Vision-Language Models (VLMs) and Large Language Models (LLMs) have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches [12, 35, 36, 39].
+For instance, the Vision-Language Frontier Maps (VLFM) method [36] leverages a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration.
+Other methods like ApexNav [39] and SG-Nav [35] build open-vocabulary representations to find novel categories.
+However, these systems inherently operate under the premise of naive instruction following—where the “target” is explicitly named (e.g., “go to the L-shape sofa”).
+They lack the complex logic required for Reasoning-based VLN needed in abstract and multi-target tasks where explicit targets are not provided (e.g., inferring that “rain” requires an “rain jacket” instead of a random coat).
+Furthermore, executing heavy foundation models all the time during the navigation process bottlenecks real-time operation on resource-constrained platforms.
 III Methodology
+Figure 2: System pipeline overview.Complex tasks are decomposed into subtasks via VLM reasoning, guided by the unified memory system.
+The first subtask always serves as input to the NeSy exploration system, which either guides exploration or sets a direct goal point for robot navigation.
+To address the limitations of existing methods in resolving abstract and multi-target tasks, we introduce VL-Nav, a neuro-symbolic vision-language navigation system that intertwines neural semantic understanding with symbolic precision.
+As shown in Fig.
+3, the architecture consists:
+  1. 1.
+    NeSy Task Planner logically decomposes complex, abstract instructions into atomic subtasks, exploration and go to, by guiding a VLM’s neural reasoning with a unified symbolic memory comprising a 3D scene graph and an object-centric image memory for task decomposition, replanning, and completion determination.
+  2. 2.
+    NeSy Exploration System couples neural semantic cues from lightweight open-vocabulary detection models (OVDMs) with a symbolic heuristic function, enabling efficient task-guided exploration of unknown environments without the computational bottlenecks of pure VLM-in-the-loop approaches.
+III-A NeSy Task Planner
+III-A1 Unified Memory System
+The Unified Memory System consists of a 3D scene graph and an object-centric image memory.
+First, we construct a 3D scene graph comprising two types of nodes: object nodes and room nodes, inspired by Hydra [16].
+Room node masks are generated using our room segmentation algorithm, which relies on morphological operations.
+To label these room nodes, we utilize a Large Language Model (LLM) that infers labels based on the object nodes connected to each room node.
+The edges between object and room nodes are established through spatial inclusion; specifically, an edge is created if an object’s centroid falls within a room’s segmentation mask.
+Object nodes are generated by an open-vocabulary detector.
+Each object node stores four key attributes: (1) the object centroid; (2) the highest-confidence detection score; (3) the robot’s pose at the viewpoint of this detection; and (4) the corresponding best-viewpoint RGB image.
+This integration of the 3D scene graph with object-centric best-viewpoint image memory forms a unified symbolic memory, providing the spatial and relational grounding necessary for subsequent task decomposition and replanning.
+III-A2 Task Decomposition and Replanning
+Leveraging Qwen3-VL [1] as the VLM backbone, the planner decomposes complex task descriptions into atomic “exploration” and “go to” subtasks.
+The VLM maintains a dynamic target object list, querying the memory module to track and register potential candidates.
+Upon the completion of each subtask, a replanning phase will be triggered, generating updated subtasks based on the updated symbolic memory.
+For target acquisition, the system employs a coarse-to-fine verification strategy.
+For instance, when locating a specific item like a “rain jacket,” the “go to” subtask is determined via a two-step process: (1) Symbolic Filtering: The 3D scene graph filters the available object nodes to propose the top-$k$ (e.g., top-3) candidates based on detection confidence.
+(2) Neural Verification: The VLM performs fine-grained reasoning on the saved best-view images and the neighbor scene graph nodes of these top candidates to determine which object best semantically matches the abstract instruction.
+Once the target is chosen, the planner publishes the saved best-view pose as the navigation goal, directing the exploration system to navigate to and report the identified target.
+Figure 3: An overview of the NeSy exploration system.
+It processes inputs including prompts, RGB images, poses, and LiDAR scans.
+The Vision-Language (VL) module conducts open-vocabulary detection to identify objects related to the prompt, generating instance-based target points.
+Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map.
+Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool.
+Finally, this system employs the NeSy scoring to select the most effective goal point.
 Figure 4: A brief illustration of VL Scoring.
-The pixel-wise open-vocabulary detection results are transferred into the spatial distibution via the Gaussian mixture model regularized by the FOV weighting (the gray arrows in the figure).
+The pixel-wise open-vocabulary detection results are transferred into the spatial distibution via the Gaussian mixture model regularized by the FoV weighting (the gray arrows).
 Then the frontier-based and the instance-based target points will be assigned with VL score based on the distribution Equation (3).
-This section presents our VL-Nav pipeline (Fig.
-2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and curiosity-vision-language (CVL) Scoring Policy to select and publish a navigation goal.
-This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
-III-A Rolling Occupancy Map: Occupancy Update
-We represent the environment as a 2D occupancy grid $\mathcal{G}$, where each cell may be free $(0)$, unknown $(-1)$, or occupied $(100)$.
-Upon receiving new sensor data (a merged obstacle-and-terrain cloud $\mathcal{P}$), we update $\mathcal{G}$ using the following steps:
-  1. 1.
-    Expand if needed. If any newly observed points lie outside the current map boundaries, expand the grid in place to preserve historical data.
-  2. 2.
-    Clear stale obstacles in forward FOV.
-For each cell marked occupied within the robot’s forward field of view (FOV), check whether it still corresponds to a point in $\mathcal{P}$.
-Any “stale” obstacle cells are re-labeled as free.
-  3. 3.
-    Inflate new obstacles.
-For each point in $\mathcal{P}$ within the sensor range $R$, mark the corresponding cell as occupied, along with a local neighborhood determined by an inflation radius $r_{\mathrm{inflation}}$.
-  4. 4.
-    Raycast unknown cells to free.
-Perform line-of-sight raycasting from the robot’s pose $(x_{r},y_{r},\theta_{r})$ across its angular range $\theta\in[\theta_{r}-\tfrac{\mathrm{hfov}}{2},\;\theta_{r}+\tfrac{\mathrm{hfov}}{2}]$.
-Any unknown cells $(-1)$ intersected by a ray are marked free $(0)$.
-Algorithm 1 summarizes these main steps, returning the updated grid $\mathcal{G}$. This occupancy update leverages consistent line-of-sight logic and inflation to support safe navigation.
-Algorithm 1 Rolling Occupancy Map Update (Key Steps)
-0:  Grid $\mathcal{G}$, merged point cloud $\mathcal{P}$, robot pose $(x_{r},y_{r},\theta_{r})$, inflation radius $r_{\mathrm{inflation}}$, half-FOV $\mathrm{hfov}/2$, sensor range $R$.
-1:  $\textsc{ExpandIfNeeded}(\mathcal{G},\mathcal{P})$
-[…]
-[…]
-[…]
-5:  return  Updated grid $\mathcal{G}$
-III-B Frontier-based Target Points
+III-B NeSy Exploration System
+Taking the subtask from the task planner, the NeSy Exploration System operationalizes high-level directives by fusing neural semantic cues from open-vocabulary detectors with symbolic geometric heuristics.
+The pipeline as shown in Fig.
+3 proceeds from the generation of candidate goal points to the hybrid NeSy scoring policy, facilitating rapid, task-guided exploration and navigation in real-time.
+III-B1 Frontier-based Target Points
 A cell $(m_{x},m_{y})$ is a frontier if it is free, $\mathcal{G}(m_{x},m_{y})=0$, and at least one neighbor is unknown, $\mathcal{G}(m_{x}+\Delta x,m_{y}+\Delta y)=-1$ for some $(\Delta x,\Delta y)$.
-After every grid update, only cells in the forward wedge are tested:
-$\begin{split}\Bigl{|}\mathrm{Angle}\bigl{(}(m_{x},m_{y}),(x_{r},y_{r})\bigr{)}-\theta_{r}\Bigr{|}\;&\leq\;\frac{\mathrm{hfov}}{2},\\ \text{and}\quad\|(m_{x},m_{y})-(x_{r},y_{r})\|\;&\leq\;R.\end{split}$
+After every grid update, only cells in the forward FoV wedge are tested:
+$\begin{split}\Bigl|\mathrm{Angle}\bigl((m_{x},m_{y}),(x_{r},y_{r})\bigr)-\theta_{r}\Bigr|\;&\leq\;\frac{\mathrm{hfov}}{2},\\ \text{and}\quad\|(m_{x},m_{y})-(x_{r},y_{r})\|\;&\leq\;R,\end{split}$
+where $(x_{r},y_{r})$ and $\theta_{r}$ denote the robot’s current position and heading, $\mathrm{hfov}$ represents the horizontal field of view, and $R$ is the maximum effective sensor range.
 A breadth-first search (BFS) clusters these frontier cells.
 Each cluster is represented by either a single centroid (for small clusters) or multiple sampled points (for large clusters) as the frontier-based target points.
-III-C Instance-Based Target Points (IBTP)
-A vision-language detector periodically reports candidate instance centers in the form $(q_{x},\;q_{y},\;\text{confidence})$, where $q_{x},q_{y}$ denote the estimated global coordinates for a potential target instance, and confidence quantifies how likely this detection is to match the desired instance.
-If confidence is larger than the detection threshold $\text{confidence is larger than the detection threshold }\tau_{\mathrm{det}}$, the candidate is retained; otherwise, it is discarded as too uncertain.
+III-B2 Instance-Based Target Points (IBTP)
+The vision-language detector shown in Fig.
+3 VL Module periodically reports candidate instance centers in the form $(q_{x},\;q_{y},\;\text{confidence})$, where $q_{x},q_{y}$ denote the estimated global coordinates for a potential target instance, and confidence quantifies how likely this detection is to match the desired instance.
+If $\text{confidence is larger than the detection threshold }\tau_{\mathrm{det}}$, the candidate is retained; otherwise, it is discarded as too uncertain.
 Retained points will be downsampled via a voxel-grid filter if many candidates lie in close proximity.
@@
 Instead, any candidate above a confidence threshold is treated as a valid goal candidate, allowing the robot to approach and verify whether it truly is an instance of interest.
-If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues, producing an accurate and robust strategy for zero-shot navigation toward a target instance.
-III-D CVL Scoring Policy
-Once frontier centroids and instance-based targets are gathered, the system computes an CVL score for each candidate goal $\mathbf{g}$.
+If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues.
+III-B3 NeSy Scoring Policy
+Once frontier-based and instance-based target points are gathered as candidate goals, the system computes the NeSy score for each candidate goal $\mathbf{g}$.
 Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance.
-The angular offset is
+The angular offset is:
 $\Delta\theta\;=\;\mathrm{Angle}\bigl(\mathbf{g},(x_{r},y_{r})\bigr)\;-\;\theta_{r}.$
@@
 As shown in Fig.
-3, a vision-language (VL) score $S_{\mathrm{VL}}(\mathbf{g})$ translates pixel-wise vision-language features into a Gaussian-mixture distribution over the robot’s horizontal field of view (FOV).
+4, a vision-language (VL) score $S_{\mathrm{VL}}(\mathbf{g})$ translates fine-grained vision-language features from the YOLO-World [6] and FastSAM [40] models into a Gaussian-mixture distribution over the robot’s horizontal field of view (FoV).
 Suppose the open-vocabulary detection model identifies $K$ likely directions or detection regions, each parameterized by $\bigl(\mu_{k},\sigma_{k},\alpha_{k}\bigr)$.
 Here, $\mu_{k}$ indicates the mean offset angle within the FoV, $\sigma_{k}$ encodes the detection’s angular uncertainty (in our implementation, a fixed constant of $0.1$ for all detections), and $\alpha_{k}$ is a confidence-based weight that captures how important that region is.
-By accumulating each region’s contribution in a mixture model, $S_{\mathrm{VL}}(\mathbf{g})$ naturally biases the robot toward goals in the direction of high-confidence detections that match the language prompt.
+By accumulating each region’s contribution in a mixture model, $S_{\mathrm{VL}}(\mathbf{g})$ naturally biases the robot toward goals in the direction of high-confidence detections that match the subtask prompt.
 From a human perspective, this process resembles observing different areas that related to the target instance with varying levels of certainty.
@@
 Inspired by VLFM [36], we then multiply by $C(\Delta\theta)$, a “view confidence” term that downweights detections with large angular offsets from the central field of view.
-The VL score is computed as:
+The VL score is computed as following equations:
 $S_{\mathrm{VL}}(\mathbf{g})\;=\;\sum_{k=1}^{K}\alpha_{k}\;\exp\!\Bigl(-\tfrac{1}{2}\Bigl(\tfrac{\Delta\theta-\mu_{k}}{\sigma_{k}}\Bigr)^{2}\Bigr)\;\cdot\;C\bigl(\Delta\theta\bigr),$
 $C(\Delta\theta)=\cos^{2}\!\biggl(\frac{\Delta\theta}{(\theta_{\mathrm{fov}}/2)}\cdot\frac{\pi}{2}\biggr).$
+Finally, to ensure the metric remains bounded, the accumulated score $S_{\mathrm{VL}}(\mathbf{g})$ is clipped to the range $[0,1]$.
 Curiosity Cues:
 We add two curiosity terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
-(1) Distance Weighting. We define
+(1) Distance Weighting. We define:
 $S_{\mathrm{dist}}(\mathbf{g})\;=\;\frac{1}{1+d(\mathbf{x}_{r},\mathbf{g})},$
@@
 (2) Unknown-Area Weighting. We further encourage curiosity-driven exploration by measuring how many unknown cells lie around $\mathbf{g}$. First, we define the ratio of unknown as:
-[…]
+$\mathrm{ratio}(\mathbf{g})\;=\;\frac{\#(\text{unknown cells})}{\#(\text{total visited cells})},$
 where a local BFS radiates outward from $\mathbf{g}$, counting how many cells are unknown versus how many are reachable (i.e., not blocked by obstacles).
@@
 To translate this raw ratio into a normalized weighting score, we apply an exponential mapping:
-[…]
+$S_{\mathrm{unknown}}(\mathbf{g})\;=\;1\;-\;\exp\!\Bigl(-\,k\,\mathrm{ratio}(\mathbf{g})\Bigr),$
 where $k$ is an adjustable parameter that controls how rapidly the score increases from 0 toward 1.
-Combined CVL Score:
+Combined NeSy Score:
 The final scores of frontier-based goals combine these three components:
-$S_{\mathrm{CVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
+$S_{\mathrm{NeSy}}(\mathbf{g})\;=\;\Bigl(w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr),$
 where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights.
-Note that the VL score is assigned when a point is initially detected, whereas the curiosity terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated.
+Note that the VL score is assigned when a point is initially detected, whereas the curiosity terms (distance and unknown-area weighting) are evaluated at goal selection time, when the current distance and unknown ratio can be calculated.
 For instance-based goals, these curiosity cues are omitted because their primary purpose is verification rather than exploration.
-III-E Select & Publish Goal Point
-After evaluating CVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate.
-Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
-Algorithm 2 Select and Publish Goal Point
-0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $CVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
-1:  $\textit{bestScore}\leftarrow-\infty$
-[…]
-3:  for all instances $\mathbf{o}_{i}$ do
-4:     if $\|\mathbf{o}_{i}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
-[…]
-6:        if $s>\textit{bestScore}$ then
-7:           $\textit{bestScore}\leftarrow s$
-8:           $\textit{bestGoal}\leftarrow\mathbf{o}_{i}$
-9:        end if
-10:     end if
-11:  end for
-[…]
-13:     for all frontiers $\mathbf{f}_{j}$ do
-14:        if $\|\mathbf{f}_{j}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
-[…]
-16:           if $s>\textit{bestScore}$ then
-17:              $\textit{bestScore}\leftarrow s$
-18:              $\textit{bestGoal}\leftarrow\mathbf{f}_{j}$
-19:           end if
-20:        end if
-21:     end for
-22:  end if
-[…]
-24:     Publish bestGoal to /selected_goal_point
-25:  end if
-26:  return
-If no candidate exceeds the distance threshold $\delta_{\mathrm{reached}}$, the robot withholds any new goal.
-Once a goal is published, a global planner updates obstacle polygons and replans a safe route.
-A local planner refines each global waypoint into short-horizon commands for real-time obstacle avoidance, supporting incremental exploration in unknown environments.
-III-F Path Planning with FAR Planner
-Once an CVL goal is selected, we hand off point-goal path planning to the FAR Planner [36], which represents obstacles polygonally and updates a visibility graph in real time.
-This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods.
-A local planner refines FAR Planner’s waypoints into short-horizon velocity commands, ensuring swift reactions to new obstacles and seamless integration with VL-Nav’s zero-shot vision-language objectives.
+III-C Goal Selection & Path Planning
+After computing scores, the system selects the optimal goal point by prioritizing high-confidence semantic matches over general exploration.
+Specifically, we iterate through all available instance-based target points that exceed a minimum distance threshold $\delta_{\mathrm{reached}}$.
+If any valid instance targets exist, the one with the highest $S_{\mathrm{VL}}$ is selected immediately to drive verification.
+If no valid instance targets are available (or all are too close), the system falls back to frontier-based target points.
+The frontier with the highest combined $S_{\mathrm{NeSy}}$ score is chosen to maximize information gain.
+If no candidate exceeds $\delta_{\mathrm{reached}}$, the robot holds its position.
+Once a goal is selected, we employ the FAR Planner [34] to generate a collision-free path, guiding the robot to the target while dynamically avoiding real-time obstacles.
 IV Experiments
-Figure 4: Four different real-world experiment environments.
-(a) Office
-(b) Apartment
-(c) Outdoor
-(d) Hallway
-Figure 5: Top-down view of the trajectories comparison on the value maps with the detection results across the four different environments.
-TABLE I: Vision-language navigation performance in 4 unseen environments (SR and SPL).
-Method SR (%) SPL
+IV-A Simulation Experiments: DARPA TIAMAT Challenge
+We first validate VL-Nav in the high-fidelity simulation environment of the DARPA TIAMAT Challenge (Phase 1).
+This benchmark is designed to test robotic reasoning under highly abstract and complex human instructions.
+Baselines:
+We compare VL-Nav against four well-known baselines and two ablation variants: closest frontier exploration, VLFM [36], SG-Nav [35], ApexNav [39], VL-Nav w/o IBTP (ablation), and VL-Nav w/o Curiosity (ablation).
+Environments & Simulators:
+The evaluation covers four distinct scenarios simulated across two platforms:
+  • HabitatSim (Indoor): The Apartment 1 and Apartment 2 environments are rendered in HabitatSim [29] to evaluate navigation in daily residential settings.
+  • IsaacSim (Outdoor): The Camping Site and Factory environments are simulated in IsaacSim [23] to leverage its advanced physics engine for unstructured settings.
+Task Settings:
+Except for interaction, each environment consists of 8 abstract navigation tasks requiring multi-step reasoning.
+To ensure reliability, we conduct 3 trials for all tasks in each environment.
+Crucially, some tasks require the robot to locate multiple target objects distributed across widely separated positions within a large-scale map.
+Instructions involve implicit semantics (e.g., “it is raining” $\rightarrow$ find umbrella), hierarchical searching (e.g., “check the garage for tools”), and abstract social concepts (e.g., “prepare for a black-tie party” $\rightarrow$ find suit, shoes, and tie).
+Robot and Sensor Setup: The DARPA TIAMAT tasks require a Boston Dynamics Spot robot, equipped with five RGB-D cameras positioned to provide a comprehensive surround view for visual reasoning.
+TABLE I: Simulation experiment results on the navigation tasks of DARPA TIAMAT challenge phase 1 in 4 different scenarios with metrics Success Rate (SR) and Max Time Usage Ratio (MTUR).
+Method SR % ($\uparrow$) MTUR ($\downarrow$)
+Apartment 1 Apartment 2 Camping Site Factory Apartment 1 Apartment 2 Camping Site Factory
+Frontier Exploration 8.3 8.3 0.0 0.0 0.958 0.884 1.000 1.000
+VLFM [36] 8.3 8.3 4.2 8.3 0.931 0.864 0.953 0.859
+SG-Nav [35] 0.0 4.2 0.0 8.3 1.000 0.973 1.000 0.901
+ApexNav [39] 25.0 25.0 20.8 12.5 0.817 0.795 0.828 0.861
+VL-Nav w/o IBTP 70.8 62.5 62.5 58.3 0.680 0.724 0.731 0.762
+VL-Nav w/o Curiosity 79.1 75.0 58.3 66.7 0.612 0.635 0.793 0.735
+VL-Nav 87.5 79.2 75.0 75.0 0.562 0.591 0.647 0.679
+Figure 5: Qualitative results in the DARPA TIAMAT Indoor-2 environment.
+We demonstrate the system’s capability to resolve three distinct types of complex instructions: (1) implicit semantic inference (weather $\rightarrow$ rain jacket); (2) hierarchical task decomposition (garage $\rightarrow$ tools); and (3) abstract concept grounding (black-tie party $\rightarrow$ formal suit).
+Figure 6: Qualitative results in the DARPA TIAMAT outdoor factory environment.
+We demonstrate robust navigation under unstructured conditions across three scenarios: (1) single-object retrieval; (2) implicit inference (measure the perimeter of a pipe $\rightarrow$ measuring tape); and (3) a challenging multi-objective task where the robot must locate three distinct targets scattered across different regions of the large-scale map.
+The colored point clouds in the insets visualize the value map guiding the exploration.
+Figure 7: Top: Four simulation environments in the DARTPA TIAMAT Challenge. Bottom: Four real-world experiment environments.
+TABLE II: Vision-language navigation performance in 4 unseen environments (SR and SPL).
+Method SR % ($\uparrow$) SPL ($\uparrow$)
 Hallway Office Apartment Outdoor Hallway Office Apartment Outdoor
@@
 VL-Nav 86.7 91.7 88.9 77.8 0.672 0.812 0.733 0.637
-IV-A Experimental Setting
-We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [39], (3) VLNav without instance-based target points, (4) VLNav without curiosity terms, and (5) the full VLNav configuration.
-Because the original VLFM relies on BLIP-2 [15], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM.
+IV-B Real-World Experiments
+Baselines:
+Since SG-Nav [35] and ApexNav [39] cannot be deployed on the edge device, we evaluate our approach in real-robot experiments against five methods: (1) closet frontier-based exploration, (2) VLFM [36], (3) VL-Nav without instance-based target points, (4) VL-Nav without curiosity terms, and (5) the full VL-Nav configuration.
+Because the original VLFM relies on BLIP-2 [17], which is too computationally heavy for real-time edge deployment, we use the same YOLO-World [6] and FastSAM [40] models instead to generate per-observation similarity scores for VLFM.
 Each method is tested under the same conditions to ensure a fair comparison of performance.
@@
 We consider four distinct environments (shown in Fig.
-4), each with a specific combination of semantic complexity (High, Medium, or Low) and size (Big, Mid, or Small).
+7), each with a specific combination of semantic complexity (High, Medium, or Low) and size (Big, Mid, or Small).
 Concretely, we use a Hallway (Medium & Big), an Office (High & Mid), an Outdoor area (Low & Big), and an Apartment (High & Small).
-In each environment, we evaluate five methods using three language prompts, yielding a diverse range of spatial layouts and semantic challenges.
-This setup provides a rigorous assessment of each method’s adaptability.
-Language-Described Instance:
-We define nine distinct, uncommon human-described instances to serve as target objects or persons during navigation.
-Examples include phrases such as “tall white trash bin,” “there seems to be a man in white,” “find a man in gray,” “there seems to be a black chair,” “tall white board,” and “there seems to be a fold chair.” The variety in these descriptions ensures that the robot must rely on vision-language understanding to accurately locate these targets.
-Robots and Sensor Setup: All experiments are conducted using a four-wheel Rover equipped with a Livox Mid-360 LiDAR.
-The LiDAR is tilted by approximately 23 degrees to the front to achieve a ±$\pm 30$ degrees vertical FOV coverage closely aligned with the forward camera’s view.
-An Intel RealSense D455 RGB-D camera, tilted upward by 7 degrees to detect taller objects, provides visual observation, though its depth data are not used for positioning or mapping.
-LiDAR measurements are a primary source of mapping and localization due to their higher accuracy.
-The whole VL-Nav system runs on an NVIDIA Jetson Orin NX on-board computer.
-IV-B Main Results
-We validate the proposed VL-Nav system in real-robot experiments across four distinct environments (Hallway, Office, Apartment, and Outdoor), each featuring different semantic levels and sizes.
-Building on the motivation articulated in Section I, we focus on evaluating VL-Nav’s ability to (1) interpret fine-grained vision-language features and conduct robust VLN, (2) explore efficiently in unfamiliar spaces across various environments, and (3) run in real-time on resource-constrained platforms.
-Fig.
-5 presents a top-down comparison of trajectories and detection results on the value map.
-Figure 6: Plots of performance in different environments sizes and semantic comlexities.
-Overall Performance:
-As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments.
-In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating CVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
-Effect of Instance-Based Target Points (IBTP):
-We note a marked improvement when enabling IBTP: the variant without IBTP lags behind, particularly in complex domains like the Apartment and Office.
-As discussed in Section III, IBTP allows VL-Nav to pursue and verify tentative detections with confidence above a threshold, mirroring human search behavior.
-This pragmatic mechanism prevents ignoring possible matches to the target description and reduces overall travel distance to confirm or discard candidate objects.
-Curiosity Contributions:
-The curiosity Score is also significant to VL-Nav’s performance. It merges two key components:
-  • Distance Weighting: Preventing easily select very far way goals to reduce travel time and energy consumption which is extremely important for the efficiency (metrics SPL) in the large-size environments.
-  • Unknown-Area Weighting: Rewards navigation toward regions that yield more information.
-Our ablations reveal that removing the distance-scoring element (VL-Nav w/o curiosity) degrades both SR and SPL, particularly in the more cluttered environments.
-Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of CVL addresses a complementary aspect of semantic navigation.
-Comparison to VLFM:
-Although the VLFM approach [39] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and CVL-based spatial reasoning.
-Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the curiosity cues via the CVL spatial reasoning.
-These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our CVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
-Summary of Findings:
-In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments.
-By fusing frontier-based target points detection, instance-based target points, and the CVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency.
-The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
-V Discussion
-V-A Sizes of Environments vs. SR and SPL
-Our experiments grouped environments by size into three categories: Small (Apartment), Mid (Office), and Big (averaging Hallway and Outdoor).
-As shown in Table I and the line plots in Fig.
-6, classical Frontier Exploration suffers a marked performance decline in larger spaces, frequently exceeding the path length limit.
-Its success rate drops from about 55.6% in the Small environment to only 36.7% in the Big, and its SPL falls from 0.564 to 0.214.
-This illustrates how naive 360° frontier-based approaches grow increasingly inefficient in larger, more open environments.
-By contrast, VL-Nav remains robust across all sizes.
-Although there is a slight decrease when moving from Mid (91.7% SR, 0.812 SPL) to Big (82.3% SR, 0.655 SPL), its overall performance remains substantially higher than the baselines.
-A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o curiosity), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
-These findings underscore two key factors.
-First, CVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig.
-1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments.
-Second, instance-based target points (IBTP) enable more targeted navigation, sustaining high success rates even as map size grows.
-V-B Semantic Complexities of Environments vs. SR and SPL
-We also vary the environments by semantic complexity, classifying them as Low (Outdoor), Medium (Hallway), and High (averaging Office and Apartment).
-An interesting trend emerges: all methods show better performance in more semantically rich environments.
-The likely explanation is that structured indoor spaces (e.g., with walls or furniture) provide stronger cues for the detection and segmentation models used by all five methods, leading to more accurate navigation.
-These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context.
-Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and curiosity-based exploration .
-The fusion of the CVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
-V-C Dynamic Environment Handling
-Handling dynamic environments is crucial for robust navigation in real-world scenarios, where changes such as moving obstacles and varying terrain can impact robot performance.
-VL-Nav incorporates several strategies to dynamically adapt to environmental changes, ensuring reliability and efficiency.
-To handle the dynamic environment, we maintain a dynamic occupancy grid map that integrates dynamic terrain analysis and path planning within the FAR planner [36].
-This allows the system to continuously update its understanding of the environment and adjust navigation strategies accordingly.
-The dynamic grid map ensures that real-time changes, such as shifting obstacles and terrain conditions, are accurately reflected, providing a more robust foundation for dynamic vision-language navigation in the social environment.
-V-D Real-Robot Efficiency Optimization
-Although VL-Nav is designed as a general navigation framework, we introduce several optimizations to ensure real-world feasibility and high performance on mobile robots.
-Visual-Language Module Variants:
-The first optimization involves selecting an efficient model variant for the visual-language (or object-detection) pipeline.
-In our tests, we ran multiple versions of YOLO-World [6] detectors with different input dimensions and TensorRT precision modes.
-For instance, one configuration used a $256\times 320$ input size in a standard GPU runtime, another used TensorRT (FP32) at the same resolution, and a third tested TensorRT (FP16) at $480\times 640$.
-Empirically, the original $256\times 320$ configuration (run without TensorRT) achieved about $23\%$ higher success rates than higher-resolution FP16 modes.
-This suggests that moderate input resolutions can strike a more favorable balance between detection accuracy and real-time performance, which is crucial for robust and efficient continuous on-robot inference.
-Rolling Occupancy Grid vs. Fixed Large Grid:
-A second optimization lies in how the occupancy grid is maintained.
-Rather than keeping a large, fixed-size global grid (common in many simulation environments), we implement a rolling grid on the real robot, dynamically shifting or expanding it only as new sensor data demand.
-This rolling approach keeps memory usage proportional to the active exploration area, lowering both storage overhead and BFS/cluster computations for partial frontier detection.
-Moreover, since real-world robots typically explore incrementally, discarding far-off map regions has minimal impact on decision-making.
-Additional System-Level Improvements:
-Finally, we incorporate various practical measures to ensure robust navigation under real-world constraints.
-For instance, the robot’s local planner and velocity commands are monitored to detect “stuck” situations; if the robot remains motionless beyond a time threshold, the system discards its current goal and reselects a new one.
-Similarly, sensor frequency and map update rates are tuned so that occupancy-grid expansions do not saturate available CPU/GPU resources.
-These measures, alongside the rolling-grid and efficient detection models, help keep VL-Nav lightweight and reactive.
-VI Limitation and Conclusion
-VI-A Limitations
-Despite the promising results demonstrated by VL-Nav, several limitations remain.
-One major challenge is the system’s ability to handle complex language descriptions.
-It struggles with spatial descriptions that contain hidden object references and objects described with specific textual annotations, which can affect navigation accuracy.
-Another limitation lies in the reliance on manually defined thresholds for various navigation conditions like the lightning condition.
-These thresholds may not generalize well across different environments and scenarios, requiring further investigation into adaptive or learning-based approaches to threshold tuning.
-VI-B Conclusion
-In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms.
-By combining pixel-wise vision-language features with a curiosity-based exploration strategy, our CVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings.
-Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%.
-The key to these gains lies in effectively leveraging semantic cues from vision-language embeddings to prioritize frontiers and potential object instances, thereby aligning navigation decisions more closely with human-like reasoning.
-Future work will explore extending VL-Nav to handle more complex instructions that involve multi-step tasks, temporal reasoning (e.g., tracking a moving target), and dynamic environments.
-Additionally, further integration with large language models could enable more nuanced command parsing and open-vocabulary object detection for even broader applicability.
-We believe these directions will help advance the capability of VLN systems and bring us closer to deploying versatile, robust robotic assistants in real-world settings.
+Task Settings:
+We define nine distinct, uncommon human-described instances to serve as target objects or persons during navigation (e.g., “tall white trash bin,” “there seems to be a man in white”).
+This results in 9 navigation tasks in total, and we conduct 3 to 5 trials for each task to robustly evaluate performance.
+This ensures that the robot must rely on NeSy reasoning to accurately locate these targets.
+Real Robot Implementation:
+Experiments are conducted on a four-wheeled Rover, with a Unitree Go2 quadruped used for the multi-floor demonstration.
+Both robots are equipped with a Livox Mid-360 LiDAR (tilted $23^{\circ}$) and an Intel RealSense D455 camera (tilted $7^{\circ}$).
+The Go2 is additionally equipped with a Unitree D1 arm and a RealSense D435 camera to enable object grasping.
+Computation is distributed: the real-time navigation stack runs onboard on an NVIDIA Jetson Orin NX, while the NeSy Task Planner runs on a remote laptop with an RTX 4090 GPU to host the Qwen3-VL-8B model [1].
+This separation is effective as the task planner operates asynchronously, updating only when a new task is received or a subtask is completed.
+IV-C Main Results
+We validate the proposed VL-Nav system through rigorous testing in both the high-fidelity DARPA TIAMAT simulation and diverse real-world deployments.
+Our evaluation employs distinct metrics suited for each domain: for simulation, we report Success Rate (SR) and Max Time Usage Ratio (MTUR), defined as the average execution time divided by the maximum time limit per task, to quantify time efficiency and completion reliability.
+For real-world experiments, we adopt SR and Success weighted by Path Length (SPL) following the VLFM convention [36] to evaluate efficiency.
+Simulation Results (DARPA TIAMAT)
+As reported in Table I, VL-Nav achieves dominant performance with SRs ranging from 75.0% to 87.5%, whereas baselines struggle.
+  • Baseline Failures and High MTUR: Standard approaches fail to resolve the abstract logic and multi-target sequencing required by the benchmark.
+Notably, SG-Nav [35] exhibits an MTUR near 1.0 across environments, frequently timing out.
+This is primarily because its reliance on LLM-based Chain-of-Thought reasoning at every navigation step incurs prohibitive latency [35].
+Similarly, ApexNav [39] suffers from low success (SR $\approx$ 12.5%–25.0%) as it lacks the consistent symbolic memory to efficiently manage long-horizon and multiple goals located at different positions.
+  • Qualitative Reasoning: The system’s ability to ground abstract concepts is visualized in Figures 5 and 6.
+In indoor settings (Fig.
+5), VL-Nav successfully infers implicit semantics (e.g., “weather” $\rightarrow$ “rain jacket”) and performs hierarchical decomposition.
+In unstructured outdoor settings (Fig.
+6), the projected value maps demonstrate how the system targets large landmarks to guide the search for smaller, distributed objects (laptop, outfit, truck) without getting lost in open space.
+Real-World Results
+Physical deployments (Table II) confirm the system’s robustness and efficiency.
+  • Efficiency (SPL): VL-Nav achieves significantly higher SPL scores (e.g., 0.812 in Office) compared to Frontier Exploration (0.317) and VLFM (0.556).
+This metric confirms that VL-Nav does not merely find targets by chance but navigates via efficient, non-redundant paths, verifying that the neuro-symbolic planning translates effectively to real-world dynamics.
+  • Sim-to-Real Transfer: The consistent performance gap between VL-Nav and baselines across both simulation and real-world experiments highlights the system’s strong generalization capabilities.
+Ablation Analysis
+  • Impact of IBTP (Verification): The w/o IBTP variant suffers most in semantically complex and cluttered environments, such as the Apartment (Sim: 87.5% $\rightarrow$ 70.8%; Real: 88.9% $\rightarrow$ 70.2%).
+Without IBTP, the robot lacks the “short-cut” mechanism to verify tentative detections, leading to missed targets and lower SPL.
+  • Impact of Curiosity (Exploration): The w/o Curiosity variant degrades primarily in large-scale, open environments, such as the Factory and Outdoor scenes (Sim Factory: 75.0% $\rightarrow$ 66.7%; Real Outdoor: 77.8% $\rightarrow$ 55.6%).
+This confirms that distance and unknown-area weighting are critical for preventing local minima and aimless wandering in expansive maps.
+Summary
+The combined results demonstrate that while baselines like SG-Nav and ApexNav stumble on the computational and logical demands of abstract multi-target tasks, VL-Nav’s neuro-symbolic architecture provides the necessary reasoning depth and execution efficiency to succeed in both high-fidelity simulation and real-world environments.
+V Conclusions and Future Work
+In this paper, we presented VL-Nav, a neuro-symbolic framework that bridges the gap between abstract human instructions and robotic execution.
+By intertwining a NeSy Task Planner with a NeSy Exploration System, our approach effectively addresses the logical and spatial reasoning challenges inherent in large-scale vision-language navigation.
+Extensive validation across the DARPA TIAMAT challenge and diverse real-world experiments demonstrates the system’s robustness, achieving overall success rates of 79.2% and 86.3%, respectively.
+These results confirm VL-Nav’s capability to complete complex tasks in large-scale unseen environments.
+Future work will focus on extending the symbolic memory to support temporal reasoning, enabling the robot to track moving targets and adapt to dynamic environmental changes.
+Additionally, we aim to incorporate lifelong learning capabilities, allowing the system to adapt to evolving environments over long-term deployments.
 Acknowledgments